OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (4, -6) and is parallel to y = -1/2x + 5.

OpenStudy (hoblos):

parallel lines have same slope so slope=m=-1/2 y+6=-1/2(x-4) simplify to get y = -1/2x -4

OpenStudy (anonymous):

oh :D ok thanks you a lot!

OpenStudy (hoblos):

my pleasure XD